# Resonance in black hole ringdown: Benchmarking quasinormal mode excitation and extraction

### v1

@@ II.4 Waveform
We address this gap and benchmark the iterative QNM extraction by considering a Kerr ringdown waveform generated by localized initial data, modeled as a delta-function source term [4]. In this special case, the QNM amplitudes are given solely by the excitation factors. Combined with the cross-validated excitation factors from the RD, our analysis enables a direct and reliable comparison of the time-domain ringdown waveform with the frequency-domain resonant excitation.
Hereafter, we focus on a contribution from single $(l,m)$ multipole. Let us consider the gravitational waveform generated by a localized source. Specifically, we adopt the following delta-function source term [4]
$T_{lm}(r^{\prime})=-r^{2}e^{-2i\omega r_{*}}\delta(r^{\prime}-r).$ (22)
@@
Figure 2: Time-domain ringdown waveforms of $(l,m)=(2,2)$ mode induced by a localized source.
On the other hand, we can analytically calculate the same integral (24) by using the residue theorem, yielding the following expression:  [^3] Note that the clockwise integration contour introduces a negative sign in the first term.
@@ III.1 Power-law tail
While the ringdown contains a superposition of an infinite number of power-law tails, we focus on the leading contribution. In principle, we could fit the tail using a fully non-linear fitting procedure. However, such an approach is often numerically unstable. To address this, we reduce the problem by considering a single-tail term
@@
characterized by amplitude $A$, power $p$, and starting time $t_{0}$. Taking the logarithm, we have
@@
This reduces the problem to a linear regression for $\log A$ and $p$ for a given non-linear parameter $t_{0}$.
@@ III.2.2 Iterative fit
Next, we determine the amplitude of the next-longest-lived mode, $C_{-0}$. We subtract the previously extracted contribution, $\Psi(t)-C_{+0}e^{-i\omega_{+0}(t)}$, fit the residual waveform using the QNM set in the second row of Table 2. We then extract $C_{-0}$, which is now the effective longest-lived mode in the residual, by the same flatness criterion, adopting the start time $t_{\text{i}}$ that minimizes $\gamma_{-0}$. This process is then iteratively repeated for subsequent modes, basically following the order of their lifetimes.
@@ IV.2 Residual waveform and contamination
Figure 6: Residual waveforms at step 8 in Figs. 4 and 5, compared with the fifth and sixth QNMs and their sum for $(l,m)=(2,2)$ at $a/M=0.9$.
@@ IV.3 Extracted amplitudes and excitation factors
Figure 8 shows the extracted amplitudes at the fitting start times at which the relative error is minimized. Although we do not yet know how to determine such optimal start times in practice, it is useful to regard them as a benchmark of the “best fitting.” In this case, the extracted amplitudes almost lie on the red line. The anticipated knot-shaped structure for the fifth and sixth overtones is still absent, but their extracted amplitudes nevertheless display a rough pattern of resonant amplification. A clearer extraction of resonance structure may require improvements to the fitting algorithm and/or the use of waveforms with smaller tail contributions.
@@ IV.4 Sharp resonances
The stable behavior of the sum is consistent with the fact that the excitation factors follow the lemniscate trajectory in the complex plane [10]. By virtue of the symmetric lemniscate pattern, the sum of the excitation factors during resonance corresponds to the center of the figure-eight shape, and hence does not change significantly. While the contribution to gravitational-wave strain scales as $\sim\omega^{-2}B$, since $\omega$ does not vary appreciably during the avoided crossing, the above argument applies, especially for sharp avoided crossings and resonances. The same conclusion holds when introducing an additional parameter controlling the sharpness of the avoided crossing and resonance, as demonstrated in Fig. S1 of Ref. [10].
@@ V Conclusion
We identified and characterized how resonances near exceptional points manifest in Kerr ringdown waveforms. We applied the iterative fitting method with mirror modes, supplemented by iterative subtraction of exact QNMs as a diagnostic check, to a ringdown gravitational waveform induced by a localized source term (24). A key advantage of employing this particular waveform is that its QNM amplitudes are directly given by the excitation factors, which have been recently computed and cross-validated. This controlled setup allowed us to investigate both mild and sharp resonances and to clarify their distinct imprints on the ringdown signal.
As a Reference Dataset (RD) of QNM frequencies and excitation factors, we used the one provided recently in Ref. [10, 11]. Complementing the three previously established methods, we have provided an additional independent validation of the RD using the confluent Heun function (see Table 1), thereby reinforcing its reliability and precision. Employing these excitation factors, we demonstrated a direct and consistent agreement between the ringdown waveform generated by localized initial data and a superposition of QNM damped sinusoids. This consistency establishes a solid basis for further studies of QNM excitations.
@@
For the mild resonance in $(l,m)=(2,2)$, we successfully extracted up to five QNM amplitudes. We achieved the extraction of the fundamental mode and the first three overtones of the original modes, and the fundamental mirror mode, all with an accuracy within 10%. The robustness of our extraction method is demonstrated by the presence of stable plateau regions, as shown in Fig. 3. For higher overtones beyond the fourth, although stable plateau regions are still observed, the extracted amplitudes deviate by more than 1010%, even though their overall magnitudes remain consistent with excitation factors. These deviations arise from the power-law tail and contamination due to error propagation from the extraction of lower overtones. As a result, our current setup is effectively limited to accurately extracting amplitudes up to around the third overtone. To enable extraction of higher overtones, more reliable fitting algorithms and/or waveforms with weaker tail contributions will be required.
@@ Appendix A Rewriting mirror modes
In the main text, we take into account the prograde and mirror modes together. The mirror modes are often rewritten in terms of the prograde mode, but it seems there exists some confusion in the literature. In this appendix, we provide a dictionary of the rewriting process.
@@
$\eqqcolon\psi_{\mathrm{originary}}+\psi_{\mathrm{mirror}}.$ (36)
As explained in the main text, we denote the prograde mode by +n+n and the mirror mode by n-n. For $m>0$, they coincide with the positive-frequency and negative-frequency modes, respectively. Since we have the relation (18) between the prograde and mirror QNM frequency, $\omega_{lm-n}=-\omega^{*}_{l-mn}$, we can rewrite $\psi_{\mathrm{retro}}$ as the sum over the prograde QNMs as
@@
$\psi_{\mathrm{originary}}+\psi_{\mathrm{mirror}}$ $=\sum_{n=+0}^{+\infty}\left[C_{lmn}e^{-i\omega_{lmn}t}S_{lm}(\theta;a\omega_{lmn})e^{im\phi}+C^{\prime}_{lmn}e^{i\omega^{*}_{l-mn}t}\{S_{l-m}(\pi-\theta;a\omega_{l-mn})e^{i(-m)\phi}\}^{*}\right].$ (41)
The final expression is consistent with expressions in the literature [see, e.g., Refs. [41, 44, 62]]. [^5] It appears that $\omega_{lmn}$ in the second term in Eq. (2.4) in Ref. [43] should be replaced by $\omega_{l-mn}$, which is consistent with the present expression (41).

### v2

@@ II.4 Waveform
We address this gap and benchmark the iterative QNM extraction by considering a Kerr ringdown waveform generated by localized initial data, modeled as a delta-function source term [33, 7]. The ringdown waveform is obtained by integration in the frequency domain [see Eq. (24) below], rather than evolving the initial data. In this special case, the QNM amplitudes are given solely by the excitation factors. Combined with the crossvalidated excitation factors from the RD, our analysis enables a direct and reliable comparison of the time-domain ringdown waveform with the frequency-domain resonant excitation.
Hereafter, we focus on a contribution from single $(l,m)$ multipole. Let us consider the gravitational waveform generated by a localized source. Specifically, we adopt the following delta-function source term [33, 7]
$T_{lm}(r^{\prime})=\frac{r^{2}e^{-2i\omega r_{*}}}{2\pi}\delta(r^{\prime}-r).$ (22)
@@
Figure 3: Time-domain ringdown waveforms of $(l,m)=(2,2)$ mode induced by a localized source. The waveform can be found in Ref. [67].
On the other hand, we can analytically calculate the same integral (24) using the residue theorem, yielding the following expression:  [^3] Note that the clockwise integration contour introduces a negative sign in the first term.
@@ III.1 Power-law tail
While the ringdown contains a superposition of an infinite number of power-law tails, we focus on the leading contribution. In principle, we could fit the tail using a fully nonlinear fitting procedure. However, in general, nonlinear fitting is often numerically unstable compared to linear fitting. To address this, we reduce the problem to considering a single-tail term
@@
characterized by complex amplitude $A$, power-law index $p$, and starting time $t_{0}$. Taking the logarithm, we have
@@
This reduces the problem to a linear regression for $\log A$ and $p$ for a given nonlinear parameter $t_{0}$.
@@ III.2.2 Iterative fit
Next, we determine the amplitude of the next-longest-lived mode, $C_{-0}$. We subtract the previously extracted contribution, $\Psi(t)-C_{+0}e^{-i\omega_{+0}(t)}$, and fit the residual waveform using the QNM set in the second row of Table 2. We then extract $C_{-0}$, which is now the effective longest-lived mode in the residual, by the same flatness criterion, adopting the start time $t_{\text{i}}$ that minimizes $\gamma_{-0}$. This process is then iteratively repeated for subsequent modes, basically following the order of their lifetimes.
@@
Note that each step carries numerical uncertainties, which propagate to subsequent steps. These numerical uncertainties contaminate the residual waveform and may eventually affect the extracted amplitudes, in particular those of higher overtones. Such contamination should be taken into account in addition to the stability of the extraction procedure, which we also address explicitly below.
@@ IV.2 Residual waveform and contamination
Figure 7: Residual waveforms at step 8 in Figs. 5 and 6, compared with the fifth and sixth QNMs and their sum for $(l,m)=(2,2)$ at $a/M=0.9$.
@@ IV.3 Extracted amplitudes and excitation factors
Figure 9 shows the extracted amplitudes at the fitting start times at which the relative error is minimized. The relative error is $\mathcal{O}(10^{-5})$ for $n=+0$, but increases to $\mathcal{O}(10^{-1})$ for $n=+5$ and $n=+6$.
Although we do not yet know how to determine such optimal start times in practice, it is useful to regard them as a benchmark of the “best fitting.” In this case, the extracted amplitudes almost lie on the red line. The anticipated knot-shaped structure for the fifth and sixth overtones is still absent, but their extracted amplitudes nevertheless display a rough pattern of resonant amplification. A clearer extraction of resonance structure may require improvements to the fitting algorithm and/or the use of waveforms with smaller tail contributions.
@@ IV.4 Sharp resonances
The stable behavior of the sum is consistent with the fact that the excitation factors follow the lemniscate trajectory in the complex plane [13]. By virtue of the symmetric lemniscate pattern, the sum of the excitation factors during resonance corresponds to the center of the figure-eight shape, and hence does not change significantly. While the contribution to gravitational-wave strain scales as $\sim\omega^{-2}B$, since $\omega$ does not vary appreciably during the avoided crossing, the above argument applies, especially for sharp avoided crossings and resonances. The same conclusion holds when introducing an additional parameter controlling the sharpness of the avoided crossing and resonance, as demonstrated in Fig. S1 of Ref. [13]. This behavior is partially consistent with Ref. [18].
@@ V Conclusion
We identified and characterized how resonances near exceptional points manifest in Kerr ringdown waveforms. We applied the iterative fitting method with mirror modes, supplemented by iterative subtraction of exact QNMs as a diagnostic check, to a ringdown gravitational waveform induced by a localized source term (24). A key advantage of employing this particular waveform is that its QNM amplitudes are directly given by the excitation factors, which have been recently computed and crossvalidated. This controlled setup allowed us to investigate both mild and sharp resonances and to clarify their distinct imprints on the ringdown signal.
As a reference dataset (RD) of QNM frequencies and excitation factors, we used the one provided recently [13, 14]. Complementing the three previously established methods, we have provided an additional independent validation of the RD using the confluent Heun function (see Table 1), thereby reinforcing its reliability and precision. Employing these excitation factors, we demonstrated a direct and consistent agreement between the ringdown waveform generated by localized initial data and a superposition of QNM damped sinusoids. This consistency establishes a solid basis for further studies of QNM excitations.
@@
For the mild resonance in $(l,m)=(2,2)$, we successfully extracted up to five QNM amplitudes. We achieved the extraction of the fundamental mode and the first three overtones of the ordinary modes, and the fundamental mirror mode, all with an accuracy within 10%. The robustness of our extraction method is demonstrated by the presence of stable plateau regions, as shown in Fig. 4. For higher overtones beyond the fourth, although stable plateau regions are still observed, the extracted amplitudes deviate by more than 1010%, even though their overall magnitudes remain consistent with excitation factors. These deviations arise from the power-law tail and contamination due to error propagation from the extraction of lower overtones. As a result, our current setup is effectively limited to accurately extracting amplitudes up to around the third overtone. To enable extraction of higher overtones, more reliable fitting algorithms and/or waveforms with weaker tail contributions will be required.
@@ Acknowledgements.
data availability
The data that support the findings of this article are openly available [14, 67, 51].
@@ Appendix A Rewriting mirror modes
In the main text, we take into account the ordinary and mirror modes together. The mirror modes are often rewritten in terms of the ordinary mode, but it seems there exists some confusion in the literature. In this appendix, we provide a dictionary of the rewriting process.
@@
$\eqqcolon\psi_{\mathrm{ordinary}}+\psi_{\mathrm{mirror}}.$ (36)
As explained in the main text, we denote the ordinary mode by +n+n and the mirror mode by n-n. For $m>0$, they coincide with the positive-frequency and negative-frequency modes, respectively. Since we have the relation (18) between the ordinary and mirror QNM frequency, $\omega_{lm-n}=-\omega^{*}_{l-mn}$, we can rewrite $\psi_{\mathrm{retro}}$ as the sum over the ordinary QNMs as
@@
$\psi_{\mathrm{ordinary}}+\psi_{\mathrm{mirror}}$ $=\sum_{n=+0}^{+\infty}\left[C_{lmn}e^{-i\omega_{lmn}t}S_{lm}(\theta;a\omega_{lmn})e^{im\phi}+C^{\prime}_{lmn}e^{i\omega^{*}_{l-mn}t}\{S_{l-m}(\pi-\theta;a\omega_{l-mn})e^{i(-m)\phi}\}^{*}\right].$ (41)
The final expression is consistent with expressions in the literature (see, e.g., Refs. [47, 50, 72]). [^5] It appears that $\omega_{lmn}$ in the second term in Eq. (2.4) in Ref. [49] should be replaced by $\omega_{l-mn}$, which is consistent with the present expression (41).